OpenStudy (anonymous):

A cylindrical tank has a radius of 3m and a depth of 10m. It is being filled at a rate of 5m^3/min. How fast is the surface rising?

OpenStudy (anonymous):

\[ V=\pi r^2h\\ \frac{dV}{dt}=\pi r^2\frac{dh}{dt}\\ 5=9\pi\frac{dh}{dt}\\ \frac{dh}{dt}=\frac5{9\pi} \]